If you wouldn't mind, could you please answer Question 1 and the corresponding subparts. Thank you so much, I really appreciate it!

Suppose you are a delivery driver and you know the relationship between the average tips you will get from one order and the distance you have to drive is:

???? = ? + ?. ? ∗ ?????

a) Draw the graph of the function above and be careful to label all variables

b) If your manager asks you to deliver one order and you know you have to drive 10 miles for this order, how much tips do you expect to get?

c) What is the marginal effect of miles on your average tips? How to interpret the marginal effect?
Answer #1

1.5 miles tips = 4 + (6) tips I (6) (c) mile tips = 4+ 1.5 (10) - 19 (Expected tip) d tips - = 1.5 d miles It says that if yo